# Singular Excitement Beyond the Horizon of a Rotating Black Hole

### v1

@@ Abstract
Previous studies have shown that an Unruh-DeWitt (UDW) detector, when coupled linearly to a massless scalar field and permitted to fall radially into certain black holes, will exhibit non-monotonicity in its transition properties near the horizon. Specifically, the transition probability of a detector freely falling into a (3+1)-dimensional Schawrzschild black hole, when considering the Unruh and Hartle-Hawking vacuum states, was shown to possess a local extremum at horizon crossing [K.K. Ng et al., New J. Phys. 24 (2022) 103018]. The transition rate of a detector falling into a static (2+1)-dimensional Bañados-Teitelboim-Zanelli (BTZ) black hole, for the Hartle-Hawking state, was also found to have multiple local extrema near the horizon under certain parameter settings [M.R. Preciado-Rivas et al., arXiv:2402.14908v1]. These discoveries are of interest, as they suggest that the event horizon of a black hole may be distinguishable to a local probe when QFT effects are included. In this paper, we explore the problem of a UDW detector falling freely into a rotating BTZ black hole. We numerically compute the detector’s transition rate for different values of black hole mass, black hole angular momentum, detector energy gap, and field boundary conditions at infinity. Our results lead to a more generalized description of the behaviour of particle detectors in BTZ black hole spacetime, from which the previous non-rotating BTZ case can be retrieved in the limit as angular momentum vanishes.
@@ I Introduction
In the absence of a quantum theory of gravity, relativistic quantum information provides valuable technical and conceptual insight for investigating quantum effects in curved spacetime. One common operational approach in RQI is to use an idealized particle detector to probe quantum fields, particularly in situations where the notion of a particle is ill-defined. The simplest such detector is the Unruh-DeWitt (UDW) detector [1, 2], a two-level quantum system that couples locally to a quantum scalar field. One quantity of interest in this framework is the transition probability between the two levels of the detector; another is the derivative of this probability with respect to total detection time, known as the transition rate. Some well-known results obtained using the UDW detector model are the Unruh effect [1, 2, 3], wherein an accelerated detector in flat spacetime exhibits a thermal response proportional to its acceleration, and entanglement degradation [4], which occurs when two detectors are in relative non-inertial motion.
The particle detector model has also been used to study black hole spacetimes. Some results obtained in this vein are described in [5], for the static detector outside of a Schwarzschild black hole, which was found to respond thermally to the Hartle-Hawking-Israel state [6]; and in [7], for a stationary detector co-rotating with a Bañados-Teitelboim-Zanelli (BTZ) black hole, which was found to thermalize at the Hawking temperature.
Much less is known about the response of detectors that freely fall into a black hole and cross its horizon. The general expectation is that the transition rate will be smooth and the detector will not thermalize. It has been shown that a detector freely falling across the horizon of a Bertotti-Robinson spacetime has a transition rate that linearly decreases as the horizon is crossed [8]. Several studies have been carried out thus far in lower-dimensional settings. [9, 10] numerically compute the transition rate of a UDW detector freely falling into a (1+1)-dimensional Schwarzschild black hole, finding evidence that the thermal properties of the detector are gradually lost during infall. In [11], the transition rate of detector infalling towards a (1+1)-dimensional Cauchy horizon was shown to diverge for a field in both the Unruh and Hartle-Hawking-Israel vacuum states. Another study [12] has shown that it is possible for two infalling detectors to harvest entanglement in (1+1)-dimensional Schwarzschild spacetime, even when causally separated by the black hole horizon. The response rate of a detector falling radially in a (2+1)-dimensional static BTZ spacetime was computed in [7], but the detector had only been switched on outside of the horizon.
@@
The transition rate of a freely falling detector in (2+1)-dimensional non-rotating BTZ spacetime was computed in [14], exhibiting similar, but richer, behaviour compared to the Schwarzschild case. In particular, the transition rate as a function of detector proper time was shown to oscillate slowly outside of the black hole, but fluctuate more dramatically near the horizon for certain ranges of the black hole’s mass and the detector’s initial position. The transition rate then increases and diverges at the singularity. Moreover, the transition rate was found to be non-differentiable at certain detector proper times; these cusps in the transition rate were dubbed “glitches.” Theoretically, one could use these glitches as an early warning system of whether the observer is about to fall into the horizon. Both the profile of the detector transition rate and the locations of the glitches are highly dependent on the black hole mass and the detector’s initial position.
@@
    We anticipate the possibility of new behaviours in detector transition rate when rotation is considered, particularly when the black hole approaches extremality. This intuition is informed by previous research showing that entanglement harvesting can be amplified by black hole rotation [15], especially as angular momentum increases to the extremal limit.
@@
We find that the detector transition rate curves for the rotating BTZ spacetime take on a similar shape as in the non-rotating BTZ case, but the rotation of the black hole pushes the features in the curves further away from the singularity. Strikingly, however, the transition rate of the detector does not diverge at the time the detector reaches the black hole’s singularity when rotation is present. Additionally, new glitch behaviours arise in the rotating case. In the non-rotating BTZ case, the detector encounters only one set of glitches, which are symmetric in the image sum. “Symmetric” in this context means that, if a glitch occurs at a proper time $\tau_{g}$ in term $n$ of the image sum, it will also occur at the same time $\tau_{g}$ in term $-n$. In the rotating BTZ spacetime, by contrast, the detector encounters four types of glitches—two of which are symmetric in $n$ and two of which are asymmetric. The sets of glitches present different behaviours for small $|n|$, but all asymptote monotonically towards the time to the inner horizon (or a reflection of this about the time to the singularity) as $n\rightarrow\pm\infty$. One can see how the special case of the non-rotating BTZ spacetime may be recovered from the rotating BTZ setting: in the limit of vanishing black hole angular momentum, the time to the inner horizon and its reflection go to the singularity; thus the glitches are collapsed into the singularity as $n\to\pm\infty$. The pattern of the glitches in the transition rate provides us with a predictor for the structure of the spacetime since it depends on the black hole mass and the detector’s switch-on moment, allowing us to anticipate if the detector will encounter the black hole horizons.
The outline of our paper is as follows. In section II, we present the basic formalism needed to compute the transition rate of a UDW detector freely falling in a rotating BTZ spacetime. We present in section III the results of our study, showing how the transition rate depends on the mass of the black hole, the angular momentum of the black hole, the boundary conditions of the field at asymptotic infinity, and the energy gap of the detector. We also present the locations of the four types of glitches in the rotating case, as well as the transition rate behaviour at the time the detector reaches the black hole singularity. In section IV, we draw conclusions from the results and propose directions for future research.
@@ II Setup
The UDW detector that we consider is a pointlike qubit, having two states denoted by $|0\rangle$ and $|E\rangle$. The energy eigenvalues of the states are 00 and $E$ respectively, where $E$ may be positive or negative. The detector starts out in the |$\ket{0}$ state, and a positive (negative) $E$ indicates that |$\ket{0}$ is the detector’s ground (excited) state. The detector moves on a timelike worldline $\mathsf{x}(\tau)$, parameterized by its proper time $\tau$, and interacts with a massless scalar field $\hat{\phi}$ via the interaction Hamiltonian
@@
$M$ is the (dimensionless) mass of the black hole, and $J$ is its angular momentum. Additionally, $t\in\mathbb{R}$, $r\in(0,\infty)$, and $\varphi\in[0,2\pi)$. This metric (4) is a vacuum solution to the Einstein field equations with cosmological constant $\Lambda=-1/\ell^{2}$, where $\ell~{}(>0)$ is the AdS length. The outer and inner horizons (where $h(r)=0$) are given by the expressions
@@
By setting $M=1$ (without loss of generality) and $\varphi\rightarrow y\in(-\infty,\infty)$ in (4), one obtains AdS3-Rindler spacetime. Conversely, the BTZ spacetime can be obtained from AdS3-Rindler by identifying $y\rightarrow\phi\in(0,\sqrt{M})$, and then rescaling $\phi$ and the radial and time coordinates to obtain (4); it is therefore locally equivalent to anti-de Sitter spacetime. This fact will be important when one considers a quantum field in BTZ spacetime, for which the correlation functions can be written as a sum of correlators in AdS3.
We pause to comment on the nature of the BTZ spacetime at $r=0$. In both the rotating and non-rotating cases, the curvature of the BTZ spacetime is constant, with a well-defined Kretschmann scalar at every point, including as $r\to 0$. However the identification that produces the BTZ spacetime is performed along a spacelike Killing vector, which is simply $\partial_{\phi}$ for the static BTZ black hole and a function of $\partial_{\phi}$ and $\partial_{t}$ for the rotating BTZ black hole.
These Killing vectors can become null and then timelike in some regions, which must be excised from the covering anti-de Sitter space to make the identifications valid; otherwise the spacetime will have closed timelike curves [22]. This in turn has the consequence that the spacetime becomes geodesically incomplete and the surface $r=0$ becomes a singularity in the causal structure [22, 23]. If $J=0$, the manifold is additionally non-Hausdorff at $r=0$ [22]. If the black hole is formed from collapsing matter then a genunine curvature singularity appears at $r=0$ [24].
We consider a freely falling detector that is initially co-rotating with the black hole. The timelike geodesic followed by the detector is given in BTZ coordinates by
@@
where $r_{0}$ is the radial position at which the detector is switched on. To simplify the notation, we will write $\tau/\ell$ as $\tau$.
@@
where $\Box$ is the d’Alembert operator, and $R$ is the Ricci scalar. The Wightman function $W(\mathsf{x},\mathsf{x}^{\prime}):=\langle 0|\hat{\phi}(\mathsf{x})\hat{\phi}(\mathsf{x}^{\prime})|0\rangle$ is the two-point correlation function in the vacuum state $|0\rangle$ of the field. Since AdS3 spacetime is locally equivalent to BTZ spacetime, we may express the field correlations in BTZ as an image sum of the Wightman function in AdS3 spacetime. The specific form is [25, 23]:
@@
where $\Gamma:(t,r,\varphi)\mapsto(t,r,\varphi+2\pi)$. The parameter $\zeta\in\{-1,0,1\}$ specifies the boundary conditions of the field at asymptotic infinity: Neumann, transparent, or Dirichlet, respectively. The Wightman function with Neumann or Dirichlet boundary conditions corresponds to a KMS state and is analytic outside of the black hole horizon, thus characterizing a Hartle-Hawking state [26, 25]. On the other hand, the state with the transparent boundary condition does not have a clear physical meaning [26, 25].
$\sigma_{\epsilon}(\mathsf{x},\mathsf{x}^{\prime})$ is the squared geodesic separation (scaled by $\ell^{2}$) between the two points in the covering AdS3 spacetime, and the $\epsilon$ notation encodes the distributional character of the Wightman function in the limit as $\epsilon\rightarrow 0_{+}$. Let the unprimed coordinates correspond to $\mathsf{x}(\tau)$ on the detector trajectory (II) and the primed coordinates to $\mathsf{x}(\tau-s)$. The explicit form of $\sigma_{\epsilon}(\mathsf{x}(\tau),\Gamma^{n}\mathsf{x}(\tau-s))$ is then
@@
Several observations can be made about the rotating case relative to the non-rotating case described in [14].
First, from (16) when calculating ˙$\dot{\mathcal{F}}_{\tau}$, the $n\neq 0$ terms are not invariant under $n\rightarrow-n$. This is different from the non-rotating case, where there was invariance in $n\rightarrow-n$. Thus, the $n$ and $-n$ terms must be summed separately in the rotating case.
Second, while the transition rate formula (16) is derived using expression (11), which is valid only in the black hole exterior, (16) holds over the detector’s full trajectory, even when the detector operates after entering the black hole. This follows from the existence of a global analytic chart, by analytic continuation in (16). This analytic continuation can be performed in both the rotating and non-rotating cases.
As in the non-rotating case, the transition rate decomposes as
@@
The square roots in (16) are positive for positive arguments, and they are analytically continued to negative arguments by giving $s$ a small negative imaginary part [7, 17]. This procedure is valid in both the rotating and non-rotating cases.
For certain values of $s$, the right hand side of (11) vanishes. Specifically, $\sigma_{\epsilon,n}(\tau,\tau-s)$ vanishes for
$s_{1,2}^{*}(\tau)=\tau-\arctan{\left(\frac{ac\pm b\sqrt{a^{2}+b^{2}-c^{2}}}{bc\mp a\sqrt{a^{2}+b^{2}-c^{2}}}\right)},$ (18)
where $a=\sin{(\tau)}K_{n,3}+\cos{(\tau)}K_{n,4}$, $b=\cos{(\tau)}K_{n,1}+\sin{(\tau)}K_{n,2}$ and $c=1$. Similarly, the quantity $\sigma_{\epsilon,n}(\tau,\tau-s)+2$ in the second term of the integrand vanishes for
$s_{3,4}^{*}(\tau)=\tau-\arctan{\left(\frac{df\pm e\sqrt{d^{2}+e^{2}-f^{2}}}{ef\mp d\sqrt{d^{2}+e^{2}-f^{2}}}\right)},$ (19)
with $d=\sin{(\tau)}K_{n,3}+\cos{(\tau)}K_{n,4}$, $e=\cos{(\tau)}K_{n,1}+\sin{(\tau)}K_{n,2}$ and $f=-1$. These values of $s$ yield singularities in the integrand of (16). However, since they are of the inverse square root type, they are still integrable.
When $s_{1,2,3,4}^{*}(\tau)$ is equal to one of the bounds of integration in (16) (i.e., $s_{1,2,3,4}^{*}(\tau)=0$ or $s_{1,2,3,4}^{*}(\tau)=\tau$), the transition rate function will gain a jump discontinuity in its first derivative at time $\tau$. Such non-differentiable points in the transition rate, dubbed “glitches,” were previously observed in the case of a UDW falling into a non-rotating BTZ black hole [14].
For the static black hole, the glitches encountered by the detector as it falls towards the singularity originate exclusively from the $s_{1}^{*}(\tau)=\tau$ case – that is, when $\sigma_{\epsilon,n}(\tau,\tau-s)$ vanishes at the upper bound $\tau$. The singularities $s^{*}_{1}(\tau)$ in the integrand have a geometric origin, occurring when the detector’s trajectory crosses the future light cone of the switch-on event. There are many discrete values of the proper time at which this can happen because the light cones wrap around the $\varphi$ direction.
@@
where the glitches at $\tau_{1}^{*}$ have the same origin as in the static case, $s^{*}_{1}$ being coincident with the upper bound of integration (i.e., $s^{*}_{1}(\tau)=\tau$). The glitches at $\tau_{3}^{*}$ arise from $s^{*}_{3}$ being coincident with the upper bound of integration, which is $s_{3}^{*}(\tau)=\tau$, but now correspond to the detector’s trajectory crossing a future-directed timelike curve beginning at the switch-on event. This is a new feature absent in the static case. The glitches at $\tau_{2}^{*}$ and $\tau_{4}^{*}$ arise from $s_{2}^{*}$ and $s_{4}^{*}$ being coincident with the lower bound of integration (i.e., $s_{2,4}^{*}(\tau)=0$) – another feature that is absent in the static case. Those from $\tau_{2}^{*}$ occur when the detector’s trajectory crosses the future light cone of the switch-on event, whereas those from $\tau_{4}^{*}$ occur when the detector intersects a future-directed timelike curve beginning at the switch-on event. For all four types of glitches, both the light cones and the future-directed timelike curves can wrap around the $\varphi$ direction, leading to many possible discrete glitches.

### v2

@@ Abstract
Previous studies have shown that an Unruh-DeWitt (UDW) detector, when coupled linearly to a massless scalar field and permitted to fall radially into certain black holes, will exhibit nonmonotonicity in its transition properties near the horizon. Specifically, the transition probability of a detector freely falling into a (3$+$1)-dimensional Schawrzschild black hole, when considering the Unruh and Hartle-Hawking vacuum states, was shown to possess a local extremum at horizon crossing [K. K. Ng et al., New J. Phys. 24, 103018 (2022)]. The transition rate of a detector falling into a static (2$+$1)-dimensional Bañados-Teitelboim-Zanelli (BTZ) black hole, for the Hartle-Hawking state, was also found to have multiple local extrema near the horizon under certain parameter settings [M.R. Preciado-Rivas et al., Phys. Rev. D 110, 025002 (2024)]. These discoveries are of interest, as they suggest that the event horizon of a black hole may be discerned by a local probe when quantum field theory effects are included. In this paper, we explore the problem of a UDW detector falling freely into a rotating BTZ black hole. We numerically compute the detector’s transition rate for different values of black hole mass, black hole angular momentum, detector energy gap, and field boundary conditions at infinity. Our results lead to a more generalized description of the behavior of particle detectors in BTZ black hole spacetime, from which the previous nonrotating BTZ case can be retrieved in the limit as angular momentum vanishes.
@@ I Introduction
In the absence of a quantum theory of gravity, relativistic quantum information provides valuable technical and conceptual insight for investigating quantum effects in curved spacetime. One common operational approach in relativistic quantum information is to use an idealized particle detector to probe quantum fields, particularly in situations where the notion of a particle is ill defined. The simplest such detector is the Unruh-DeWitt (UDW) detector [1, 2], a two-level quantum system that couples locally to a quantum scalar field. One quantity of interest in this framework is the transition probability between the two levels of the detector; another is the derivative of this probability with respect to total detection time, known as the transition rate. Some well-known results obtained using the UDW detector model are the Unruh effect [1, 2, 3], wherein an accelerated detector in flat spacetime exhibits a thermal response proportional to its acceleration, and entanglement degradation [4], which occurs when two detectors are in relative noninertial motion.
The particle detector model has also been used to study black hole spacetimes. Some results obtained in this vein are described in [5], for the static detector outside of a Schwarzschild black hole, which was found to respond thermally to the Hartle-Hawking-Israel state [6]; and in [7], for a stationary detector corotating with a Bañados-Teitelboim-Zanelli (BTZ) black hole, which was found to thermalize at the Hawking temperature.
Much less is known about the response of detectors that freely fall into a black hole and cross its horizon. The general expectation is that the transition rate will be smooth and the detector will not thermalize. It has been shown that a detector freely falling across the horizon of a Bertotti-Robinson spacetime has a transition rate that linearly decreases as the horizon is crossed [8]. Several studies have been carried out thus far in lower-dimensional settings. References [9, 10] numerically compute the transition rate of a UDW detector freely falling into a (1+1)-dimensional Schwarzschild black hole, finding evidence that the thermal properties of the detector are gradually lost during infall. In [11], the transition rate of a detector infalling toward a (1+1)-dimensional Cauchy horizon was shown to diverge for a field in both the Unruh and Hartle-Hawking-Israel vacuum states. Another study [12] has shown that it is possible for two infalling detectors to harvest entanglement in (1+1)-dimensional Schwarzschild spacetime, even when causally separated by the black hole horizon. The response rate of a detector falling radially in a (2+1)-dimensional static BTZ spacetime was computed in [7], but the detector had only been switched on outside of the horizon.
@@
The transition rate of a freely falling detector in (2+1)-dimensional nonrotating BTZ spacetime was computed in [14], exhibiting similar, but richer, behavior compared to the Schwarzschild case. In particular, the transition rate as a function of detector proper time was shown to oscillate slowly outside the black hole, but fluctuate more dramatically near the horizon for certain ranges of the black hole’s mass and the detector’s initial position. The transition rate then increases and diverges at the singularity. Moreover, the transition rate was found to be nondifferentiable at certain detector proper times; these cusps in the transition rate were dubbed “glitches.” Theoretically, one could use these glitches as an early warning system as to whether or not the observer is about to fall into the horizon. Both the profile of the detector transition rate and the locations of the glitches are highly dependent on the black hole mass and the detector’s initial position.
@@
    We anticipate the possibility of new behaviors in detector transition rate when rotation is considered, particularly when the black hole approaches extremality. This intuition is informed by previous research showing that entanglement harvesting can be amplified by black hole rotation [15], especially as angular momentum increases to the extremal limit.
@@
We find that the detector transition rate curves for the rotating BTZ spacetime take on a similar shape as in the nonrotating BTZ case, but the rotation of the black hole pushes the features in the curves further away from the singularity. However, the transition rate of the detector does not diverge at the time the detector reaches its point of closest approach to the black hole’s singularity when rotation is present. Additionally, new glitch behaviors arise in the rotating case. In the nonrotating BTZ case, the detector encounters only one set of glitches, which are symmetric in the image sum. “Symmetric” in this context means that, if a glitch occurs at a proper time $\tau_{g}$ in term $n$ of the image sum, it will also occur at the same time $\tau_{g}$ in term $-n$. In the rotating BTZ spacetime, by contrast, the detector encounters four types of glitches—two of which are symmetric in $n$ and two of which are asymmetric. The sets of glitches present different behaviors for small $|n|$, but all asymptote monotonically toward the time to the inner horizon (or a reflection of this about $\Delta\tau/\ell=\pi/2$) as $n\rightarrow\pm\infty$. One can see how the special case of the nonrotating BTZ spacetime may be recovered from the rotating BTZ setting: in the limit of vanishing black hole angular momentum, the time to the inner horizon and its reflection go to the singularity; thus, the glitches are collapsed into the singularity as $n\to\pm\infty$. The pattern of the glitches in the transition rate provides us with a predictor for the structure of the spacetime since it depends on the black hole mass and the detector’s switch-on moment, allowing us to anticipate if the detector will encounter the black hole horizons.
The outline of our paper is as follows. In section II, we present the basic formalism needed to compute the transition rate of a UDW detector freely falling in a rotating BTZ spacetime. We present in section III the results of our study, showing how the transition rate depends on the mass of the black hole, the angular momentum of the black hole, the boundary conditions of the field at asymptotic infinity, and the energy gap of the detector. We also present the locations of the four types of glitches in the rotating case, as well as the transition rate behavior at the time the detector reaches its closest approach to the black hole singularity. In section IV, we draw conclusions from the results and propose directions for future research.
@@ II Setup
The UDW detector that we consider is a pointlike qubit, having two states denoted by $|0\rangle$ and $|E\rangle$. The energy eigenvalues of the states are 00 and $E$, respectively, where $E$ may be positive or negative. The detector starts out in the |$\ket{0}$ state, and a positive (negative) $E$ indicates that |$\ket{0}$ is the detector’s ground (excited) state. The detector moves on a timelike worldline $\mathsf{x}(\tau)$, parametrized by its proper time $\tau$, and interacts with a massless scalar field $\hat{\phi}$ via the interaction Hamiltonian
@@
$M$ is the (dimensionless) mass of the black hole, and $J$ is its angular momentum. Additionally, $t\in\mathbb{R}$, $r\in(0,\infty)$, and $\varphi\in[0,2\pi)$. This metric (4) is a vacuum solution to the Einstein field equations with cosmological constant $\Lambda=-1/\ell^{2}$, where $\ell~{}(>0)$ is the anti-de Sitter (AdS) length. The outer and inner horizons (where $h(r)=0$) are given by the expressions
@@
By setting $M=1$ (without loss of generality) and $\varphi\rightarrow y\in(-\infty,\infty)$ in (4), one obtains the AdS3-Rindler spacetime. Conversely, the BTZ spacetime can be obtained from AdS3-Rindler by identifying $y\rightarrow\phi\in(0,\sqrt{M})$, and then rescaling $\phi$ and the radial and time coordinates to obtain (4); it is therefore locally equivalent to anti-de Sitter spacetime. We consider this fact when dealing with a quantum field in BTZ spacetime, where the correlation functions can be represented as a sum of correlators in AdS3.
We pause to comment on the nature of the BTZ spacetime at $r=0$. In both the rotating and nonrotating cases, the curvature of the BTZ spacetime is constant, with a well-defined Kretschmann scalar at every point, including as $r\to 0$. However the identification that produces the BTZ spacetime is performed along a spacelike Killing vector, which is simply $\partial_{\phi}$ for the static BTZ black hole and a function of $\partial_{\phi}$ and $\partial_{t}$ for the rotating BTZ black hole.
These Killing vectors can become null and then timelike in some regions, which must be excised from the covering anti-de Sitter space to make the identifications valid; otherwise, the spacetime will have closed timelike curves [22]. This, in turn, has the consequence that the spacetime becomes geodesically incomplete and the surface $r=0$ becomes a singularity in the causal structure [22, 23]. If $J=0$, the manifold is additionally non-Hausdorff at $r=0$ [22]. If the black hole is formed from collapsing matter then a genuine curvature singularity appears at $r=0$ [24].
We consider a freely falling detector that is initially corotating with the black hole. The timelike geodesic followed by the detector is given in BTZ coordinates by
@@
where $r_{0}$ is the radial position at which the detector is switched on, and constant terms in the temporal and angular coordinates have been omitted. To simplify the notation, we will write $\tau/\ell$ as $\tau$.
@@
where $\Box$ is the d’Alembert operator, and $R$ is the Ricci scalar. The Wightman function $W(\mathsf{x},\mathsf{x}^{\prime}):=\langle 0|\hat{\phi}(\mathsf{x})\hat{\phi}(\mathsf{x}^{\prime})|0\rangle$ is the two-point correlation function in the vacuum state $|0\rangle$ of the field. Since AdS3 spacetime is locally equivalent to BTZ spacetime, we may express the field correlations in BTZ as an image sum of the Wightman function in AdS3 spacetime. The specific form is [25, 23]
@@
where $\Gamma:(t,r,\varphi)\mapsto(t,r,\varphi+2\pi)$. The parameter $\zeta\in\{-1,0,1\}$ specifies the boundary conditions of the field at asymptotic infinity: Neumann, transparent, or Dirichlet, respectively. The Wightman function with Neumann or Dirichlet boundary conditions corresponds to a Kubo-Martin-Schwinger (KMS) state and is analytic outside the black hole horizon, thus characterizing a Hartle-Hawking state [26, 25]. On the other hand, the state with the transparent boundary condition does not have a clear physical meaning [26, 25].
$\sigma_{\epsilon}(\mathsf{x},\mathsf{x}^{\prime})$ is the squared geodesic separation (scaled by $\ell^{2}$) between two points in the covering AdS3 spacetime, and the $\epsilon$ notation encodes the distributional character of the Wightman function in the limit as $\epsilon\rightarrow 0_{+}$. Let the unprimed coordinates correspond to $\mathsf{x}(\tau)$ on the detector trajectory (II) and the primed coordinates to $\mathsf{x}(\tau-s)$. The explicit form of $\sigma_{\epsilon}(\mathsf{x}(\tau),\Gamma^{n}\mathsf{x}(\tau-s))$ is then
@@
Several observations can be made about the rotating case relative to the nonrotating case described in [14].
First, from (16) when calculating ˙$\dot{\mathcal{F}}_{\tau}$, the $n\neq 0$ terms are not invariant under $n\rightarrow-n$. This is different from the nonrotating case, where there was invariance in $n\rightarrow-n$. Thus, the $n$ and $-n$ terms must be summed separately in the rotating case.
Second, while the transition rate formula (16) is derived using expression (11), which is valid only in the black hole exterior, (16) holds over the detector’s full trajectory, even when the detector operates after entering the black hole. This follows from the existence of a global analytic chart, by analytic continuation in (16). This analytic continuation can be performed in both the rotating and nonrotating cases.
As in the nonrotating case, the transition rate decomposes as
@@
The square roots in (16) are positive for positive arguments, and they are analytically continued to negative arguments by giving $s$ a small negative imaginary part [7, 17]. This procedure is valid in both the rotating and nonrotating cases.
For certain values of $s$, the right-hand side of (11) vanishes. Specifically, $\sigma_{\epsilon,n}(\tau,\tau-s)$ vanishes for
$\begin{split}s_{1,2}^{*}(\tau)=\tau-\operatorname{arctan2}&\left({a\mp b\sqrt{a^{2}+b^{2}-1}},\right.\\ &\hskip 5.80002pt\left.{b\pm a\sqrt{a^{2}+b^{2}-1}}\right),\end{split}$ (18)
[…]
$\begin{split}s_{3,4}^{*}(\tau)=\tau-\operatorname{arctan2}&\left({-a\mp b\sqrt{a^{2}+b^{2}-1}},\right.\\ &\hskip 4.10004pt\left.{-b\pm a\sqrt{a^{2}+b^{2}-1}}\right).\end{split}$ (19)
These values of $s$ yield singularities in the integrand of (16). However, since they are of the inverse square root type, they are still integrable.
When $s_{1,2,3,4}^{*}(\tau)$ is equal to one of the bounds of integration in (16) (i.e., $s_{{\color[rgb]{0.88671875,0.265625,0.51171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.88671875,0.265625,0.51171875}1,3}}^{*}(\tau)={\color[rgb]{0.88671875,0.265625,0.51171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.88671875,0.265625,0.51171875}\tau}$ or $s_{{\color[rgb]{0.88671875,0.265625,0.51171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.88671875,0.265625,0.51171875}2,4}}^{*}(\tau)={\color[rgb]{0.88671875,0.265625,0.51171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.88671875,0.265625,0.51171875}0}$), the transition rate function will gain a jump discontinuity in its first derivative at time $\tau$. Such nondifferentiable points in the transition rate, dubbed “glitches,” were previously observed in the case of a UDW detector falling into a nonrotating BTZ black hole [14].
For the static black hole, the glitches encountered by the detector as it falls toward the singularity originate exclusively from the $s_{1}^{*}(\tau)=\tau$ case—that is, when $\sigma_{\epsilon,n}(\tau,\tau-s)$ vanishes at the upper bound $\tau$. The singularities $s^{*}_{1}(\tau)$ in the integrand have a geometric origin, occurring when the detector’s trajectory crosses the future light cone of the switch-on event. There are many discrete values of the proper time at which this can happen because the light cones wrap around the $\varphi$ direction.
@@
where the glitches at $\tau_{1}^{*}$ have the same origin as in the static case, $s^{*}_{1}$ being coincident with the upper bound of integration (i.e., $s^{*}_{1}(\tau)=\tau$). The glitches at $\tau_{3}^{*}$ arise from $s^{*}_{3}$ being coincident with the upper bound of integration, which is $s_{3}^{*}(\tau)=\tau$, but now correspond to the detector’s trajectory crossing a future-directed timelike curve beginning at the switch-on event. This is a new feature absent in the static case. The glitches at $\tau_{2}^{*}$ and $\tau_{4}^{*}$ arise from $s_{2}^{*}$ and $s_{4}^{*}$ being coincident with the lower bound of integration (i.e., $s_{2,4}^{*}(\tau)=0$)—another feature that is absent in the static case. Those from $\tau_{2}^{*}$ occur when the detector’s trajectory crosses the future light cone of the switch-on event, whereas those from $\tau_{4}^{*}$ occur when the detector intersects a future-directed timelike curve beginning at the switch-on event. For all four types of glitches, both the light cones and the future-directed timelike curves can wrap around the $\varphi$ direction, leading to many possible discrete glitches.